The recommended retail price of a brand of designer jeans is known. The price of the jeans in a sample of 16 retailers is on average $141 with a sample standard deviation of 4. If this is a ‘random’ sample and the prices can be assumed to be normally distributed, construct a 95% confidence interval for the average sale pric

1
Expert's answer
2022-01-31T16:43:16-0500

The critical value for "\alpha = 0.05" and "df = n-1 = 15" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =2.131449."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(141-2.131449\times\dfrac{4}{\sqrt{16}}, 141+2.131449\times\dfrac{4}{\sqrt{16}})"

"=(138.87, 143.13)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "138.87 < \mu < 143.137," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(138.87, 143.13)."